part of the proceeds from a garage sale was 290 worth of $5 and $20 bills, if there were 3 more $5 bills than $20 bills. find the number of each denomination
Answer by fractalier(6550)

Let x = # of $5 bills.
Let y = # of $20 bills.
The value equation is
5x + 20y = 290
The other equation is
x = y + 3
Substituting this into the first equation gives
5(y + 3) + 20y = 290
5y + 15 + 20y = 290
25y = 275
y = 11
x = 14